OpenStudy (anonymous):

(x+y)dx + (x-y)dy=0 (EXACT) answer pls.

OpenStudy (anonymous):

the equation is exact

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Mdx+Ndy=0 M=x+y N=x-y

OpenStudy (anonymous):

\[\int\limits (x+y) dx=x^2/2+xy+H(y)\]

OpenStudy (anonymous):

is that the one it seems short solution

OpenStudy (anonymous):

differntiate partially with respect to y x+H'(y)=N x+H'(y)= x-y H'(y)=-y H(y)=-y^2/2

OpenStudy (anonymous):

the solution is x^2/2+xy-y^2/2=C

OpenStudy (anonymous):

u mean that's the general SOLUTION isn't it

OpenStudy (anonymous):

how about this - y= 6e^x - 7e^2x + e^-x (4+3x)

OpenStudy (anonymous):

answer me pls.

OpenStudy (anonymous):

the sol. i posted is the gen sol what do u want?

OpenStudy (anonymous):

yeah thank you how about my nxt question.. how about this - y= 6e^x - 7e^2x + e^-x (4+3x)

OpenStudy (anonymous):

what to do with it?

OpenStudy (anonymous):

Obtain the solution

OpenStudy (anonymous):

cant get u.. its not DE

OpenStudy (anonymous):

this one means ^x = up the x

OpenStudy (anonymous):

It's an differential equations

OpenStudy (anonymous):

its an equation, where are the differentials?

OpenStudy (anonymous):

ok if it is an equation could you Obtain the Solution. ?